# via Mixture-of-Curvature Experts

### v1

@@ Abstract
Frontier large language models (LLMs) have shown great success in text modeling and generation tasks across domains. However, natural language exhibits inherent semantic hierarchies and nuanced geometric structure, which current LLMs do not capture completely owing to their reliance on Euclidean operations such as dot-products and norms. Furthermore, recent studies have shown that not respecting the underlying geometry of token embeddings leads to training instabilities and degradation of generative capabilities. These findings suggest that shifting to non-Euclidean geometries can better align language models with the underlying geometry of text. We thus propose to operate fully in Hyperbolic space, known for its expansive, scale-free, and low-distortion properties. To this end, we introduce HELM, a family of HypErbolic Large Language Models, offering a geometric rethinking of the Transformer-based LLM that addresses the representational inflexibility, missing set of necessary operations, and poor scalability of existing hyperbolic LMs. We additionally introduce a Mixture-of-Curvature Experts model, HELM-MiCE, where each expert operates in a distinct curvature space to encode more fine-grained geometric structure from text, as well as a dense model, HELM-D. For HELM-MiCE, we further develop hyperbolic Multi-Head Latent Attention (HMLA) for efficient, reduced-KV-cache training and inference. For both models, we further develop essential hyperbolic equivalents of rotary positional encodings and root mean square normalization. We are the first to train fully hyperbolic LLMs at billion-parameter scale, and evaluate them on well-known benchmarks such as MMLU and ARC, spanning STEM problem-solving, general knowledge, and commonsense reasoning. Our results show consistent gains from our HELM architectures – up to 4% – over popular Euclidean architectures used in LLaMA and DeepSeek, highlighting the efficacy and enhanced reasoning afforded by hyperbolic geometry in large-scale language model pretraining.
@@ 1 Introduction
Contemporary Large Language Models (LLMs) [18, 42, 9, 1] fundamentally operate within Euclidean space. This manifests in their reliance on Euclidean operations such as dot products and norms applied to token embeddings. However, this architecture presents a potential mismatch with the intrinsic structure of natural language data. Existing works have shown that textual data, particularly token inputs to LLMs, exhibit an inherent semantic hierarchy [48, 47, 21, 35], thus requiring a space that can naturally accommodate these relationships. An ideal LLM architecture would possess geometric alignment with the underlying structure of the data it aims to represent.
@@
To further illustrate the unique geometry of text data, we show in Figure 1 the distribution of Ricci Curvature of token embeddings from popular decoder-only LLMs on 1000 diverse samples from RedPajama [46] [^1] Ricci Curvature is a metric of how strongly or loosely connected a region of space is. The more negative the curvature, the more hyperbolic the space is. We describe the metric in Appendix A.1.. We observe that the vast majority of tokens exhibit a wide range of negative curvatures. This has also been observed by Robinson et al. [38], who investigate the subspace of token embeddings and its inherent non-Euclidean structure. As Ricci curvature measures the local geometry of a manifold, these empirical observations suggest hierarchical token structures, while the variation in curvature values suggest complex token geometry that cannot be captured by single curvature approaches. Robinson et al. [38] also show that not respecting the geometry of tokens will harm a Transformer-based LLM’s generative capabilities while introducing undue training instabilities. He et al. [21] also document power law distributions of token frequencies, implying hierarchy among tokens. We thus propose a geometric rethinking of the transformer-based LLM: operate fully in hyperbolic space, where the negative curvature results in exponentially increasing volume w.r.t. distance. Hyperbolic space is expansive, scale-free, and yields low distortion, and has shown success in numerous settings, particularly in Transformer architectures [48, 5, 19, 6, 39]. It provides token embeddings more “breathing room” in the network, better aligning with the underlying structure of the text data.
However, previous works that study hyperbolic Transformers and pre-trained models have several major shortcomings: (1) Inflexible geometric spaces. They assign each Transformer block a single hyperbolic manifold, embedding entire token sequences in fixed-curvature spaces. This approach does not align with the observed substantial variant in curvature values as noted above, thus limiting the expressiveness of hidden representations. (2) Lack of essential operations. They omit widely-used LLM components such as rotary positional encoding and RMS normalization, and lack theoretical guarantees of LLM modules in Euclidean settings; (3) Poor scalability. They focus on low-dimensional settings and use quadratic hyperbolic self-attention mechanisms that do not scale comparably to modern Euclidean foundation models [18, 10]. In this work, we address the limitations of both Euclidean LLMs and prior hyperbolic Transformers through the following contributions:
@@ 2 Related Work
Hyperbolic Transformers.  Hyperbolic geometry has emerged as a powerful embedding space for representational learning, particularly in domains characterized by hierarchical and scale-free structures [16, 34, 15]. Hyperbolic Neural Networks (HNN) [15] and their extension HNN++ [39] have shown that operating in hyperbolic spaces increases representational capacity. Several works have incorporated hyperbolic geometry into Transformers to better capture semantic relationships. HAT [19], HNN++ [39], and HyboNet [5] all propose equivalent formulations of hyperbolic attention in different models of hyperbolic space. HypFormer [48] developed several essential modules that were lacking in previous works, showing improved performance for structured and hierarchical datasets. Although work on hyperbolic pre-trained language models exist [6], they ignore essential components required to train large language models, such as normalization layers, residual connections, roatry positional encodings. Furthermore, these works suffer from the limitations mentioned in Section 1. Previous work have also considered mixed-curvature Transformers [7], however, they only using different curvature values in each attention head and relying on tangent space methods that are prone to mapping errors and numerical stability [5, 48, 20, 3]. Further related works include LResNet [23] who introduces an efficient hyperbolic residual connection method that mitigated numerical instability and tangent space mappings. Some works have devised hybrid approaches by incorporating hyperbolic components to existing pre-trained Euclidean LLMs and vision models [12, 36, 47, 17, 49, 31]. However, these works do not develop and pre-train hyperbolic LLMs from scratch. Our work advances this line of research by developing and scaling hyperbolic architectures to large-scale pretraining setups, and additionally introducing novel components for efficient and expressive language modeling in hyperbolic space.
@@
Open Large Language Models.  The recent surge in open LLMs has democratized access to cutting-edge NLP capabilities, enabling broader research and application development. Notable among these is LLaMA [44], which introduced a family of efficient and powerful models trained on diverse, large-scale corpora. Llama models employ several optimizations such as rotary positional embeddings and grouped query attention, making them competitive across various downstream tasks. Building on these ideas, Gemma [42] introduced further improvements for better data curation, advanced pertaining techniques, and careful model scaling strategies. In parallel, Mixture-of-Experts (MoE) architectures have emerged as a prominent design to enhance model capacity without a proportional increase in computation cost. DeepSeek-MoE [11] introduces an efficient routing mechanism to dynamically activate a subset of experts per input, significantly improving inference throughput compared to other MoE models such as Mixtral [27]. However, all these models are inherently Euclidean, and while effective, may not align well with the geometry of the token subspace.
@@ 3.1 Lorentz Hyperbolic Geometry
There are several isometric models of hyperbolic space employed in previous research [45, 19, 43, 34, 48, 5, 3]. In this study, we choose the Lorentz model: its special space-time interactions allow for fully hyperbolic operations, offering enhanced expressiveness and empirical stability from an optimization perspective [5, 48, 33]. Nevertheless, our methods can be easily reformulated for other models of hyperbolic geometry via isometric mappings.
Lorentz model. An $n$-dimensional Lorentz model, denoted as $\mathbb{L}^{K,n}$, is a Riemannian manifold $\mathcal{L}^{n}$ equipped with the Riemannian metric tensor $\mathfrak{g}_{n}^{K}=\mathrm{diag}(-1,1,\ldots,1)$ and defined by a constant negative curvature $K<0$. Each point $\mathbf{x}\in\mathbb{L}^{K,n}$ has a parametrized form $[x_{t},\mathbf{x}_{s}]^{T}$, where $x_{t}\in\mathbb{R}$ is the time-like dimension and $\mathbf{x}_{s}\in\mathbb{R}^{n}$ is the space-like dimension. For points $\mathbf{x},\mathbf{y}\in\mathbb{L}^{K,n}$, their Lorentzian inner product $\langle\mathbf{x},\mathbf{y}\rangle_{\mathcal{L}}$ is given by $\langle\mathbf{x},\mathbf{y}\rangle_{\mathcal{L}}=-x_{t}y_{t}+\mathbf{x}_{s}^{T}\mathbf{y}_{s}=\mathbf{x}^{T}\mathfrak{g}_{n}^{K}\mathbf{y}.$ Hence, the Lorentzian norm $|\|\mathbf{x}\||_{\mathcal{L}}\coloneq\sqrt{|\langle\mathbf{x},\mathbf{x}\rangle_{\mathcal{L}}|}$. Formally, the point set $\mathcal{L}^{n}\coloneq\{\mathbf{x}\in\mathbb{R}^{n+1}:\langle\mathbf{x},\mathbf{x}\rangle_{\mathcal{L}}=1/K,x_{t}>0\}$. The origin $\mathbf{o}\in\mathbb{L}^{K,n}$ is the point $[\sqrt{-1/K},0,\ldots,0]^{T}$.
@@ 3.2 Hyperbolic Neural Network Modules
Extensive work has been done to develop hyperbolic neural network modules [48, 5, 3, 45, 23, 15, 39, 19], which we introduce below.
Lorentz Linear Layer. Given curvatures $K_{1},K_{2}$, and parameters $\mathbf{W}\in\mathbb{R}^{(n+1)\times m}$ and $\mathbf{b}\in\mathbb{R}^{m}$, the Lorentzian linear transformation with curvature [48] is the map $\mathrm{HLT}:\mathbb{L}^{K_{1},n}\to\mathbb{L}^{K_{2},m}$ given by,
@@
Lorentz Residual Connection. Let $\mathbf{x},f(\mathbf{x})\in\mathbb{L}^{K,n}$ where $\mathbf{x}$ is an input vector and $f(\mathbf{x})$ is the output of a neural network $f$. Then, the Lorentzian residual connection [23] is given by,
@@
Hyperbolic self-attention. Hyperbolic self-attention is defined equivalently in different models through manifold midpoints [39, 19, 5]. Given $T$ tokens $\mathbf{X}$ where $\mathbf{x}_{i}\in{\mathbb{L}^{K,n}}$, $\mathbf{W^{Q}},\mathbf{W^{K}},\mathbf{W^{V}}\in\mathbb{R}^{(n+1)\times d}$, then the Lorentzian self-attention [5, 48] is given by
$\mathrm{Att}_{\mathcal{L}}(\mathbf{x})_{i}=\frac{\sum_{j=1}^{T}\nu_{i,j}\mathbf{v}_{j}}{\sqrt{-K}|\|\sum_{l=1}^{T}\nu_{i,k}\mathbf{v}_{l}\||_{\mathcal{L}}},$ where $\text{where }\nu_{i,j}=\frac{\exp\left(-d_{\mathcal{L}}^{2}(\mathbf{q}_{i},\mathbf{k}_{j})/\sqrt{m}\right)}{\sum_{l=1}^{T}\exp\left(-d^{2}_{\mathcal{L}}(\mathbf{q}_{i},\mathbf{k}_{l})/\sqrt{m}\right)}$ (3)
@@ 4.1 Hyperbolic Rotary Positional Encoding
Previous works that proposed positional encoding methods in hyperbolic space [5, 48, 22] are confined to only learning relative encodings. However, contemporary LLMs [18, 10, 4] have instead shifted towards Rotary Positional Encodings (RoPE) [40], a scalable alternative that incorporates aspects from both absolute and relative encoding methods, improving length generalization and downstream performance. We thus propose a novel hyperbolic rotary positional encoding (HoPE) to construct positional encodings in hyperbolic space, and prove the same theoretical guarantees as RoPE. Given $T$ tokens $\mathbf{X}$, where $\mathbf{x}_{i}\in\mathbb{L}^{K,d}$ ($d$ even), let $\mathbf{Q},\mathbf{K}$ be the queries and keys as in Equation 3. The hyperbolic rotary positional encoding applied to the $i$-th token is,
@@
Next, we study the theoretical aspects of HoPE; all proofs can be found in Appendix A. First note that since $\mathbf{R}_{i,\Theta}$ is an Euclidean rotation matrix, it isometrically preserves the (Euclidean) norm of vectors. Given the definition of the Lorentz model (Section 3.1), an equivalent expression for Equation 4 is $\mathrm{HoPE}(\mathbf{z}_{i})=\begin{pmatrix}1&0\\ 0&\mathbf{R}_{i,\Theta}\end{pmatrix}\mathbf{z}_{i}$, making $\mathrm{HoPE}$ a valid Lorentzian rotation operation (see, e.g., [5]).
Validity. A defining characteristic for the Euclidean RoPE is that the inner product of the encoded keys and queries is a function of only the word embeddings and their relative position. Thus, only the relative positional information is encoded [40]. For hyperbolic attention in Equation 3, the analogous is defined with $-d^{2}_{\mathcal{L}}$ instead, which we formalize below.
@@
Let $\mathbf{X}$ be $T$ tokens with $\mathbf{x}_{i}\in\mathbb{L}^{K,d}$. Let $\mathbf{Q},\mathbf{K}$ be queries and keys as in Equation 3. Then $-d^{2}_{\mathcal{L}}\left(\mathrm{\mathrm{HoPE}\left(\mathbf{q}_{a}\right),\mathrm{HoPE}\left(\mathbf{k}_{b}\right))}\right)=g(\mathbf{x}_{a},\mathbf{x}_{b};a-b)$ for some function $g$.
HoPE only encodes relative positional information based on Proposition 4.1 similar to RoPE, which establishes its validity as a RoPE operation.
@@
Let $\mathbf{Q},\mathbf{K}$ be as defined in Equation 3, then the negative square Lorentz distance $-d_{\mathcal{L}}\left(\mathrm{HoPE}\left(\mathbf{q}_{i}\right),\mathrm{HoPE}(\mathbf{k}_{j})\right)$ can be upper bounded by $f(\mathbf{q}_{i},\mathbf{k}_{j})g(i-j)<0$, where $f$ has no dependencies on position, and $g$ depends entirely on relative position and scales inversely w.r.t. $i-j$.
Thus, HoPE ensures far-apart tokens have weaker connections based on Proposition 4.2, a property not guaranteed by previous learned encoding methods.
Robustness to arbitrary token distances. Barbero et al. [2] recently show that decaying token dependency (Proposition 4.2) may not be the primary reason for RoPE’s success, but rather it enables LLMs to attend to specific relative distances. Our formulation of HoPE also ensures this property:
@@
Let $\mathbf{Q},\mathbf{K}$ be as defined in Equation 3, then HoPE can be maximal at an arbitrary distance, i.e., for any relative distance $r\in\mathbb{Z}$, there exists a key $\mathbf{k}_{j}$ such that the softmax value is maximum at distance $r$.
@@
Positional attention. Barbero et al. [2] also prove that using RoPE can help capture purely positional relationships via diagonal attention patterns, where tokens only attend to themselves, and off-diagonal attention patterns, where tokens attend only to their preceding neighbor. HoPE also allows for this:
@@
We provide ablations comparing HoPE to prior hyperbolic positional encodings in Appendix D.

### v2

@@ Abstract
Frontier large language models (LLMs) have shown great success in text modeling and generation tasks across domains. However, natural language exhibits inherent semantic hierarchies and nuanced geometric structure, which current LLMs do not capture completely owing to their reliance on Euclidean operations such as dot-products and norms. Furthermore, recent studies have shown that not respecting the underlying geometry of token embeddings leads to training instabilities and degradation of generative capabilities. These findings suggest that shifting to non-Euclidean geometries can better align language models with the underlying geometry of text. We thus propose to operate fully in Hyperbolic space, known for its expansive, scale-free, and low-distortion properties. To this end, we introduce HELM, a family of HypErbolic Large Language Models, offering a geometric rethinking of the Transformer-based LLM that addresses the representational inflexibility, missing set of necessary operations, and poor scalability of existing hyperbolic LMs. We additionally introduce a Mixture-of-Curvature Experts model, HELM-MiCE, where each expert operates in a distinct curvature space to encode more fine-grained geometric structure from text, as well as a dense model, HELM-D. For HELM-MiCE, we further develop hyperbolic Multi-Head Latent Attention (HMLA) for efficient, reduced-KV-cache training and inference. For both models, we further develop essential hyperbolic equivalents of rotary positional encodings and root mean square normalization. We are the first to train fully hyperbolic LLMs at billion-parameter scale, and evaluate them on well-known benchmarks such as MMLU and ARC, spanning STEM problem-solving, general knowledge, and commonsense reasoning. Our results show consistent gains from our HELM architectures – up to 4% – over popular Euclidean architectures used in LLaMA and DeepSeek with superior semantic hierarchy modeling capabilities, highlighting the efficacy and enhanced reasoning afforded by hyperbolic geometry in large-scale language model pretraining.
@@ 1 Introduction
Contemporary Large Language Models (LLMs) [dubey2024llama3, gemma2024, dai2024deepseekmoe, achiam2023gpt] fundamentally operate within Euclidean space. This manifests in their reliance on Euclidean operations such as dot products and norms applied to token embeddings. However, this architecture presents a potential mismatch with the intrinsic structure of natural language data. Existing works have shown that textual data, particularly token inputs to LLMs, exhibit an inherent semantic hierarchy [yang2024hypformer, yang2024hyplora, he2025position, nickel2018learning], thus requiring a space that can naturally accommodate these relationships. An ideal LLM architecture would possess geometric alignment with the underlying structure of the data it aims to represent.
@@
To further illustrate the unique geometry of text data, we show in Figure 1 the distribution of Ricci Curvature of token embeddings from popular decoder-only LLMs on 1000 diverse samples from RedPajama [weber2024redpajamaopendatasettraining] [^1] Ricci Curvature is a metric of how strongly or loosely connected a region of space is. The more negative the curvature, the more hyperbolic the space is. We describe the metric in Appendix A.1.. We observe that the vast majority of tokens exhibit a wide range of negative curvatures. This has also been observed by robinson2024structuretokenspacelarge, who investigate the subspace of token embeddings and its inherent non-Euclidean structure. As Ricci curvature measures the local geometry of a manifold, these empirical observations suggest hierarchical token structures, while the variation in curvature values suggest complex token geometry that cannot be captured by single curvature approaches. robinson2024structuretokenspacelarge also show that not respecting the geometry of tokens will harm a Transformer-based LLM’s generative capabilities while introducing undue training instabilities. he2025position also document power law distributions of token frequencies, implying hierarchy among tokens. We thus propose a geometric rethinking of the transformer-based LLM: operate fully in hyperbolic space, where the negative curvature results in exponentially increasing volume w.r.t. distance. Hyperbolic space is expansive, scale-free, and yields low distortion, and has shown success in numerous settings, particularly in Transformer architectures [yang2024hypformer, chen2021fully, gulcehre2019hyperbolicAT, chen2024hyperbolic, HNN++]. It provides token embeddings more “breathing room” in the network, better aligning with the underlying structure of the text data.
However, previous works that study hyperbolic Transformers and pre-trained models have several major shortcomings: (1) Inflexible geometric spaces. They assign each Transformer block a single hyperbolic manifold, embedding entire token sequences in fixed-curvature spaces. This approach does not align with the observed substantial variant in curvature values as noted above, thus limiting the expressiveness of hidden representations. (2) Lack of essential operations. They omit widely-used LLM components such as rotary positional encoding and RMS normalization, and lack theoretical guarantees of LLM modules in Euclidean settings; (3) Poor scalability. They focus on low-dimensional settings and use quadratic hyperbolic self-attention mechanisms that do not scale comparably to modern Euclidean foundation models [dubey2024llama3, deepseekai2024deepseekv3technicalreport]. In this work, we address the limitations of both Euclidean LLMs and prior hyperbolic Transformers through the following contributions:
@@ 2 Related Work
Hyperbolic Transformers.  Hyperbolic geometry has emerged as a powerful embedding space for representational learning, particularly in domains characterized by hierarchical and scale-free structures [ganea2018hyperbolic, nickel2017poincare, HNN]. Hyperbolic Neural Networks (HNN) [HNN] and their extension HNN++ [HNN++] have shown that operating in hyperbolic spaces increases representational capacity. Several works have incorporated hyperbolic geometry into Transformers to better capture semantic relationships. HAT [gulcehre2019hyperbolicAT], HNN++ [HNN++], and HyboNet [chen2021fully] all propose equivalent formulations of hyperbolic attention in different models of hyperbolic space. HypFormer [yang2024hypformer] developed several essential modules that were lacking in previous works, showing improved performance for structured and hierarchical datasets. Although work on hyperbolic pre-trained language models exist [chen2024hyperbolic], they ignore essential components required to train large language models, such as normalization layers, residual connections, roatry positional encodings. Furthermore, these works suffer from the limitations mentioned in Section˜1. Previous work have also considered mixed-curvature Transformers [cho2023curve], however, they only using different curvature values in each attention head and relying on tangent space methods that are prone to mapping errors and numerical stability [chen2021fully, yang2024hypformer, he2016deep, Bdeir2024fully]. Further related works include LResNet [he2025lresnet] who introduces an efficient hyperbolic residual connection method that mitigated numerical instability and tangent space mappings. Some works have devised hybrid approaches by incorporating hyperbolic components to existing pre-trained Euclidean LLMs and vision models [desai2023hyperbolic, pal2025compositional, yang2024hyplora, Ge2022HyperbolicCL, yue2023hyperbolic, mandica2024hyperbolic]. However, these works do not develop and pre-train hyperbolic LLMs from scratch. Our work advances this line of research by developing and scaling hyperbolic architectures to large-scale pretraining setups, and additionally introducing novel components for efficient and expressive language modeling in hyperbolic space.
@@
Open Large Language Models.  The recent surge in open LLMs has democratized access to cutting-edge NLP capabilities, enabling broader research and application development. Notable among these is LLaMA [touvron2023llama], which introduced a family of efficient and powerful models trained on diverse, large-scale corpora. Llama models employ several optimizations such as rotary positional embeddings and grouped query attention, making them competitive across various downstream tasks. Building on these ideas, Gemma [gemma2024] introduced further improvements for better data curation, advanced pertaining techniques, and careful model scaling strategies. In parallel, Mixture-of-Experts (MoE) architectures have emerged as a prominent design to enhance model capacity without a proportional increase in computation cost. DeepSeek-MoE [deepseekv2] introduces an efficient routing mechanism to dynamically activate a subset of experts per input, significantly improving inference throughput compared to other MoE models such as Mixtral [jiang2024mixtral]. However, all these models are inherently Euclidean, and while effective, may not align well with the geometry of the token subspace.
@@ 3.1 Lorentz Hyperbolic Geometry
There are several isometric models of hyperbolic space employed in previous research [van2023poincar, gulcehre2019hyperbolicAT, tifrea2018poincar, nickel2017poincare, yang2024hypformer, chen2021fully, Bdeir2024fully]. In this study, we choose the Lorentz model: its special space-time interactions allow for fully hyperbolic operations, offering enhanced expressiveness and empirical stability from an optimization perspective [chen2021fully, yang2024hypformer, mishne2024numericalstabilityhyperbolicrepresentation]. Nevertheless, our methods can be easily reformulated for other models of hyperbolic geometry via isometric mappings.
Lorentz model. An $n$-dimensional Lorentz model, denoted as $\mathbb{L}^{K,n}$, is a Riemannian manifold $\mathcal{L}^{n}$ equipped with the Riemannian metric tensor $\mathfrak{g}_{n}^{K}=\mathrm{diag}(-1,1,\ldots,1)$ and defined by a constant negative curvature $K<0$. Each point $\mathbf{x}\in\mathbb{L}^{K,n}$ has a parametrized form $x_{t},\mathbf{x}_{s}]^{T}$, where $x_{t}\in\mathbb{R}$ is the time-like dimension and $\mathbf{x}_{s}\in\mathbb{R}^{n}$ is the space-like dimension. For points $\mathbf{x},\mathbf{y}\in\mathbb{L}^{K,n}$, their Lorentzian inner product $\langle\mathbf{x},\mathbf{y}\rangle_{\mathcal{L}}$ is given by $\langle\mathbf{x},\mathbf{y}\rangle_{\mathcal{L}}=-x_{t}y_{t}+\mathbf{x}_{s}^{T}\mathbf{y}_{s}=\mathbf{x}^{T}\mathfrak{g}_{n}^{K}\mathbf{y}.$ Hence, the Lorentzian norm $|\|\mathbf{x}\||_{\mathcal{L}}\coloneq\sqrt{|\langle\mathbf{x},\mathbf{x}\rangle_{\mathcal{L}}|}$. Formally, the point set $\mathcal{L}^{n}\coloneq\{\mathbf{x}\in\mathbb{R}^{n+1}:\langle\mathbf{x},\mathbf{x}\rangle_{\mathcal{L}}=1/K,x_{t}>0\}$. The origin $\mathbf{o}\in\mathbb{L}^{K,n}$ is the point $\sqrt{-1/K},0,\ldots,0]^{T}$.
@@ 3.2 Hyperbolic Neural Network Modules
Extensive work has been done to develop hyperbolic neural network modules [yang2024hypformer, chen2021fully, Bdeir2024fully, van2023poincar, he2025lresnet, HNN, HNN++, gulcehre2019hyperbolicAT], which we introduce below.
Lorentz Linear Layer. Given curvatures $K_{1},K_{2}$, and parameters $\mathbf{W}\in\mathbb{R}^{(n+1)\times m}$ and $\mathbf{b}\in\mathbb{R}^{m}$, the Lorentzian linear transformation with curvature [yang2024hypformer] is the map $\mathrm{HLT}:\mathbb{L}^{K_{1},n}\to\mathbb{L}^{K_{2},m}$ given by,
@@
Lorentz Residual Connection. Let $\mathbf{x},f(\mathbf{x})\in\mathbb{L}^{K,n}$ where $\mathbf{x}$ is an input vector and $f(\mathbf{x})$ is the output of a neural network $f$. Then, the Lorentzian residual connection [he2025lresnet] is given by,
@@
Hyperbolic self-attention. Hyperbolic self-attention is defined equivalently in different models through manifold midpoints [HNN++, gulcehre2019hyperbolicAT, chen2021fully]. Given $T$ tokens $\mathbf{X}$ where $\mathbf{x}_{i}\in{\mathbb{L}^{K,n}}$, $\mathbf{W^{Q}},\mathbf{W^{K}},\mathbf{W^{V}}\in\mathbb{R}^{(n+1)\times d}$, then the Lorentzian self-attention [chen2021fully, yang2024hypformer] is given by
$\mathrm{Att}_{\mathcal{L}}(\mathbf{x})_{i}=\frac{\sum_{j=1}^{T}\nu_{i,j}\mathbf{v}_{j}}{\sqrt{-K}|\|\sum_{l=1}^{T}\nu_{i,k}\mathbf{v}_{l}\||_{\mathcal{L}}},$ $\text{where }\nu_{i,j}=\frac{\exp\left(-d_{\mathcal{L}}^{2}(\mathbf{q}_{i},\mathbf{k}_{j})/\sqrt{m}\right)}{\sum_{l=1}^{T}\exp\left(-d^{2}_{\mathcal{L}}(\mathbf{q}_{i},\mathbf{k}_{l})/\sqrt{m}\right)}$ (3)
@@ 4.1 Hyperbolic Rotary Positional Encoding
Previous works that proposed positional encoding methods in hyperbolic space [chen2021fully, yang2024hypformer, he2025hypercore] are confined to only learning relative encodings. However, contemporary LLMs [dubey2024llama3, deepseekai2024deepseekv3technicalreport, gpt-neox-20b] have instead shifted towards Rotary Positional Encodings (RoPE) [su2021roformer], a scalable alternative that incorporates aspects from both absolute and relative encoding methods, improving length generalization and downstream performance. We thus propose a novel hyperbolic rotary positional encoding (HoPE) to construct positional encodings in hyperbolic space, and prove the same theoretical guarantees as RoPE. Given $T$ tokens $\mathbf{X}$, where $\mathbf{x}_{i}\in\mathbb{L}^{K,d}$ ($d$ even), let $\mathbf{Q},\mathbf{K}$ be the queries and keys as in Equation˜3. The hyperbolic rotary positional encoding applied to the $i$-th token is,
@@
Next, we study the theoretical aspects of HoPE; all proofs can be found in Appendix˜A. First note that since $\mathbf{R}_{i,\Theta}$ is an Euclidean rotation matrix, it isometrically preserves the (Euclidean) norm of vectors. Given the definition of the Lorentz model (Section˜3.1), an equivalent expression for Equation˜4 is $\mathrm{HoPE}(\mathbf{z}_{i})=\begin{pmatrix}1&0\\ 0&\mathbf{R}_{i,\Theta}\end{pmatrix}\mathbf{z}_{i}$, making $\mathrm{HoPE}$ a valid Lorentzian rotation operation (see, e.g., [chen2021fully]).
Validity. A defining characteristic for the Euclidean RoPE is that the inner product of the encoded keys and queries is a function of only the word embeddings and their relative position. Thus, only the relative positional information is encoded [su2021roformer]. For hyperbolic attention in Equation˜3, the analogous is defined with $d^{2}_{\mathcal{L}}$ instead, which we formalize below.
@@
Let $\mathbf{X}$ be $T$ tokens with $\mathbf{x}_{i}\in\mathbb{L}^{K,d}$. Let $\mathbf{Q},\mathbf{K}$ be queries and keys as in Equation˜3. Then $d^{2}_{\mathcal{L}}\left(\mathrm{\mathrm{HoPE}\left(\mathbf{q}_{a}\right),\mathrm{HoPE}\left(\mathbf{k}_{b}\right))}\right)=g(\mathbf{x}_{a},\mathbf{x}_{b};a-b)$ for some function $g$.
HoPE only encodes relative positional information based on Proposition˜4.1 similar to RoPE, which establishes its validity as a RoPE operation.
@@
Let $\mathbf{Q},\mathbf{K}$ be as defined in Equation˜3, then the negative square Lorentz distance $d_{\mathcal{L}}\left(\mathrm{HoPE}\left(\mathbf{q}_{i}\right),\mathrm{HoPE}(\mathbf{k}_{j})\right)$ can be upper bounded by $f(\mathbf{q}_{i},\mathbf{k}_{j})g(i-j)<0$, where $f$ has no dependencies on position, and $g$ depends entirely on relative position and scales inversely w.r.t. iji-j.
Thus, HoPE ensures far-apart tokens have weaker connections based on Proposition˜4.2, a property not guaranteed by previous learned encoding methods.
Robustness to arbitrary token distances. barbero2025rope recently show that decaying token dependency (Proposition 4.2) may not be the primary reason for RoPE’s success, but rather it enables LLMs to attend to specific relative distances. Our formulation of HoPE also ensures this property:
@@
Let $\mathbf{Q},\mathbf{K}$ be as defined in Equation˜3, then HoPE can be maximal at an arbitrary distance, i.e., for any relative distance $r\in\mathbb{Z}$, there exists a key $\mathbf{k}_{j}$ such that the softmax value is maximum at distance $r$.
@@
Positional attention. barbero2025rope also prove that using RoPE can help capture purely positional relationships via diagonal attention patterns, where tokens only attend to themselves, and off-diagonal attention patterns, where tokens attend only to their preceding neighbor. HoPE also allows for this:
@@
We provide ablations comparing HoPE to prior hyperbolic positional encodings in Appendix˜D.